The lenghts of the sides of a triangle are x, 2x+2, and 13. If the perimeter of the triangle is 30, what is the value of x?
Answer by hinaahsan(5)


Perimeter of triangle = x+2x+2+13
x+2x+2+13=30
3x+15=30
3x=30-15
3x=15
x=15/3
x=5